# Timing performance of SiPM-on-Tile elements: Laboratory and test beam measurements

### v1

@@ 4.3 Results: Light Collection
Since the electronics and light detection do not contribute significantly to the overall timing properties of the signals that we observe at the test beam, the two remaining processes are light collection and scintillation. The upper panel of figure 12 shows the photon time distributions of these two processes for a BC408 $30\times 30\times 3~{}\text{mm}^{3}$ scintillator tile: The light collection was measured with the laser setup and the scintillation histogram is the same as in figure 8.
@@
The lower panel of figure 12 shows the photon time distribution for a BC408 $30\times 30\times 3~{}\text{mm}^{3}$ tile measured at the test beam. The time structure of the test beam signals agrees very well with the red line, which is a convolution of the scintillation and light collection distributions from the upper plot.
This demonstrates that the signals that we observe at the test beam are indeed determined by scintillation and light collection together. The comparison shown in figure 12 works equally well for the scintillator sizes $20\times 20\times 3~{}\text{mm}^{3}$ and $40\times 40\times 3~{}\text{mm}^{3}$, and can be found in [16].
@@
Figure 12: Top: Photoelectron time distribution of scintillation and light collection for BC408 $30\times 30\times 3~{}\text{mm}^{3}$ scintillator tiles. Bottom: Photoelectron time spectrum observed at the test beam with the same tile geometry, and a convolution of scintillation and light collection. The photon hit times were extracted by decomposition of the waveform into 1 p.e. peaks (see section 3.1).
The light collection is a geometric effect that arises from the different paths that photons travel in the scintillator tile. Figure 13 compares the photon time distribution for the light collection measurements with a laser on three different scintillator tile sizes. There is clearly a different time structure for these three geometries, with smaller tiles responding faster. This suggests that the reason why the time resolution of the SiPM-on-tile setup depends on the geometry of the scintillator tile is not only the light yield, but also the time response arising from different light paths in the scintillator.
(a) Light collection
(b) Test beam
Figure 13: Photoelectron time distribution obtained at test beam measurements (right) and from light collection only (left) for different scintillator tile sizes. The photon hit times were extracted by decomposition of the waveform into 1 p.e. peaks (see section 3.1).
@@ 5 Conclusions
for an energy deposition corresponding to the most probable value of a minimum-ionizing particle. The time resolution scales with the inverse of the square root of the signal amplitude.
In order to disentangle the contributions of scintillation, light collection and light detection to the timing properties, we have developed different measurement methods where the signals are created using electrons from a radioactive 90Sr source or pulsed laser light.
@@
The analysis of test beam data shows that the time resolution strongly depends on the energy deposition in the scintillator tile, and thus also on the light yield. The light yield in turn depends on the scintillator tile size. Moreover, the laser measurements suggest that the light collection in the scintillator tile also has a time structure that depends on the scintillator tile size.
This motivates to further study the timing properties of the SiPM-on-tile technology, and specifically how the scintillator tile size, the light yield, and the time resolution are correlated. In an upcoming publication these open questions will be addressed with a Geant4-based simulation of the SiPM-on-tile setup.
@@ Acknowledgements
We thank our colleagues in the CALICE Collaboration for helpful discussions.

### v2

@@ 5.2 Scintillation and Light Collection
Since the electronics and light detection do not contribute significantly to the overall timing properties of the signals that we observe at the test beam, the two remaining processes are light collection and scintillation. The upper panels of figure 10 show the photon time distributions of these two processes for BC408 scintillator tiles: The light collection was measured with the laser setup and the scintillation histogram is the same as in figure 7(b).
@@
The lower panels of figure 10 show the photon time distribution for BC408 scintillator tiles measured at the test beam. The time structures of the test beam signals agree very well with the red lines, which are convolutions of the scintillation and light collection distributions from the upper plots. This demonstrates that the signals that we observe at the test beam are indeed determined by scintillation and light collection together.
(a) $20\times 20\times 3~\text{mm}^{3}$
(b) $30\times 30\times 3~\text{mm}^{3}$
(c) $40\times 40\times 3~\text{mm}^{3}$
@@
Figure 10: Top row: The photoelectron time distribution of the light collection process for BC408 scintillator tiles of different sizes is shown in orange. The green histogram is the photon emission time distribution of BC408 which is also shown in figure 7(b). Bottom row: Photoelectron time spectrum observed at the test beam with the same tile geometry, and a convolution of scintillation and light collection. The photon hit times were extracted by decomposition of the waveforms into 1 p.e. peaks (see section 3).
5.3 Quantitative Discussion of Photon Time Distributions
For a quantitative discussion we calculate the rise time, fall time and FWHM of the histograms shown in figure 10 using a Monte Carlo method. The following procedure is applied to each of the histograms:
  • 10000 random photon hits are generated by sampling from the histogram.
  • The randomly generated photons are again binned into a histogram with a bin of width 0.4 ns, and a univariate spline is calculated as an interpolation function.
  • The rise time, fall time and FWHM are calculated from the univariate spline function. The rise time and fall time are defined as the time between 20 % and 80 % of the maximum signal amplitude.
  • The above process is repeated 1000 times, and a mean and standard deviation for rise time, fall time and FWHM are calculated.
The calculated statistics for the photon time distributions are visualized in figure 11.
(a) Rise Time
(b) Fall Time
(c) FWHM
Figure 11: Rise time, fall time and FWHM of photon time distributions for test beam measurements, scintillation and light collection. The values of all data points are listed in table 5.
The rise time of the light collection process shows a slight increase with tile size which indicates that in a smaller scintillator tile there is a smaller time spread of the first photons that reach the SiPM. However, the rise times of the overall process observed at the beam test show no clear trend for different tile sizes.
The fall times and FWHMs of the photon time distributions measured at the beam test show a clear dependence on the scintillator tile size. This means that larger scintillator tiles generally produce a wider photon time distribution, which can also be seen in figure 10. A possible explanation is that the photons have a longer mean free path in a larger scintillator tile which increases the time over which they can reach the SiPM. This would mean that the loss of photons is not dominated by absorption in the scintillator material [^4] According to the manufacturer, BC408 has an absorption length of 2.1m and a refractive index of 1.58. Therefore, one would expect a decay constant of approximately 11 ns for all scintillator tile sizes, if the photon losses were dominated by absorption in the scintillator., but rather by inefficient reflection at the ESR foil or the edge of the tile.
The rise time, fall time and FWHM of the light collection alone, measured with laser pulses, are smaller than the corresponding numbers inferred for the scintillation process even for the largest scintillator tiles tested. This indicates that, while the tile geometry has an impact on the time resolution, a large contribution comes from the scintillator material itself. In the SiPM-on-tile geometries tested, the time spread of the photons measured at the SiPM is therefore not limited by the geometry but by the scintillator material.
The data points from figure 11 are also listed in table 5. For reference, the data sheet values of rise and fall time of BC408 are shown together with the measurements presented in this paper. Since the manufacturer does not state how their time constants are defined, these values are however not directly comparable.
Table 5: Rise time, fall time and FWHM of photon time distributions for test beam measurements, scintillation and light collection. These values are visualized in figure 11.
Rise time (ns) Fall time (ns) FWHM (ns)
BC408 time distribution measured 0.97 $\pm$ 0.07 3.36 $\pm$ 0.14 2.70 $\pm$ 0.21
BC408 manufacturer information [15] 0.9 2.1
Light collection $20\times 20\times 3~\text{mm}^{3}$ 0.60 $\pm$ 0.04 0.55 $\pm$ 0.03 0.757 $\pm$ 0.014
Light collection $30\times 30\times 3~\text{mm}^{3}$ 0.71 $\pm$ 0.03 1.43 $\pm$ 0.11 1.11 $\pm$ 0.03
Light collection $40\times 40\times 3~\text{mm}^{3}$ 0.74 $\pm$ 0.02 1.91 $\pm$ 0.08 1.22 $\pm$ 0.05
Test Beam $20\times 20\times 3~\text{mm}^{3}$ 0.97 $\pm$ 0.09 4.26 $\pm$ 0.34 4.32 $\pm$ 0.27
Test Beam $30\times 30\times 3~\text{mm}^{3}$ 1.23 $\pm$ 0.09 5.3 $\pm$ 0.5 5.16 $\pm$ 0.20
[…]
Since the naive scaling of the time resolution with overall light yield, given in equation 2.3, implicitly assumes that the shape of time distribution of the photons is constant for different amplitudes, this description needs to be refined when considering different tile sizes. The time resolution is determined based on the rising flank of the signal, so only the fraction of the photons in the rising part of the signal, $r_{rise}$, contribute to the time resolution. We approximate that this fraction is proportional to the inverse of the width of the photon time distribution, $r_{rise}\propto 1/FWHM$. Therefore, equation 2.3 is modified by introducing an additional term that corrects for the geometry dependence of the fraction of the signal contributing to the time measurement, given by the inverse of the width of the photon time distribution, resulting in
$\sigma_{t}\propto\frac{1}{\sqrt{n_{photons}\cdot r_{rise}}}\propto\sqrt{\frac{FWHM}{LY}}\propto A^{1/4}\cdot FWHM^{1/2}.$ (5.3)
It is thus expected that the quantity $\sigma_{1MIP}\cdot A^{-1/4}\cdot FWHM^{-1/2}$ is constant for different tile sizes. The calculated values, shown in table 6, agree with this hypothesis within $2\sigma$, significantly better than with a simple scaling shown in table 3 which only takes into account the amplitude alone.
Table 6: Comparison of the scaling of the time resolution with the tile area and FWHM of the photon time distribution. The right column is expected to be constant for all tile sizes, if the the approximation in equation 5.3 holds.
Tile size $\sigma_{1MIP}\cdot A^{-1/4}\cdot FWHM^{-1/2}$ ($\sqrt{\text{ns}/\text{cm}}$)
Test Beam $20\times 20\times 3~\text{mm}^{3}$ 0.131 $\pm$ 0.008
Test Beam $30\times 30\times 3~\text{mm}^{3}$ 0.147 $\pm$ 0.009
Test Beam $40\times 40\times 3~\text{mm}^{3}$ 0.15 $\pm$ 0.01
@@ 6 Conclusions
for an energy deposition corresponding to the most probable value of a minimum-ionizing particle. The time resolution scales with the inverse of the square root of the signal amplitude. This means that the time resolution strongly depends on the energy deposition in the scintillator tile, and thus also on the light yield. The light yield in turn depends on the scintillator tile size.
In order to disentangle additional contributions of scintillation, light collection and light detection to the timing properties, we have developed different measurement methods where the signals are created using electrons from a radioactive 90Sr source or pulsed laser light.
@@
The light collection process shows a clear dependence on the scintillator tile size, with larger scintillator tiles delivering broader photon time distributions. While this contribution is small compared to the emission time of the scintillator material itself, it does result in an additional tile-size dependence of the time resolution beyond the effect of the light yield alone, with further improved time resolution for smaller scintillator tiles.
In order to quantify how the scintillator tile size, the light yield, and the time resolution are correlated, we have developed a Geant4-based simulation of a SiPM-on-Tile detector which will be presented in an upcoming publication.
@@ Acknowledgements
We thank our colleagues in the CALICE Collaboration for helpful discussions. We also express our gratitude to the anonymous referee for the excellent suggestions during the review process, which led to substantial improvements of this paper.